morganislame:

What is the equation, in point-slope form, for a line that goes through (8,−4) and has a slope of −56 ? y+4=−56(x−8) y+4=−56(x+8) y−4=−56(x+8) y−4=−56(x−8)

563blackghost:

point-slope is setup as \(\large\bf{y-y_{1}=m(x-x_{1})}\). \(\bf{m}\) is the slope. \(\bf{(x_{1},y_{1})}\) is the point given. So we plug. \(\large\bf{y+4=-56(x-8)}\)

morganislame:

Thank you so much. That is what i was thinking but was a little bit confused.

563blackghost:

np :)